OpenStudy (anonymous):

Can the expression s sqrt(3) + 27 sqrt(2^5) + 9 sqrt(3s) be simplified into an expression with fewer terms? Explain and simplify, if possible.

OpenStudy (anonymous):

I know it simplifies to s sqrt(3) + 3s^2 sqrt(3s) + 9 sqrt(3s) can it be simplified to fewer terms? How and why?

zepdrix (zepdrix):

This is the original expression? :o \[\Large s \sqrt3+27\sqrt{2^5}+9\sqrt{3s}\]

OpenStudy (anonymous):

Yes it is

zepdrix (zepdrix):

All three terms contain different degrees of s, so no we can't combine them any further. At least not without doing some fancy factoring business :d\[\Large \color{royalblue}{s^1} \sqrt3+\color{royalblue}{s^0}27\sqrt{2^5}+\color{royalblue}{s^{1/2}}9\sqrt{3}\] s^1/2 = sqrt s s^0 = 1, i just wanted to throw those in to show that they're different D:

OpenStudy (anonymous):

So what is the final answer? I'm confused, sorry

zepdrix (zepdrix):

`Can the expression be simplified into an expression with fewer terms?` No. `Explain and simplify, if possible.` Simplification not possible. I hope my explanation helped a tiny bit :C I'm not sure exactly how to explain it.

zepdrix (zepdrix):

Like do you understand why `these` three terms can't be combined together?\[\Large x^2+x+1\]They're different degrees of x, so they don't combine, yes? :o Where as something like:\[\Large x^2+2x^2+1\]Would simplify to:\[\Large 3x^2+1\]Since the squared terms can be combined.

zepdrix (zepdrix):

Your problem is the same idea, the degrees on s are just not as clearly written out.

OpenStudy (anonymous):

So once it's simplified to \[s \sqrt{3} + 27^{2}\sqrt{2} + 9\sqrt{3s}\] that's as far as it goes?

OpenStudy (anonymous):

Yeah, I now understand the degree thing!

OpenStudy (anonymous):

Wait! I know that thing was incorrect!

zepdrix (zepdrix):

ya I'm not sure what happened there :) lol your middle term kinda magically changed to something else!

OpenStudy (anonymous):

i know :/ haha okay so once I get here I can't go any further? s sqrt(3) + 3s^2 sqrt(3s) + 9 sqrt(3s) even though the two last terms have the same degree of s?

zepdrix (zepdrix):

Hmm I don't understand what's going on, your middle term keeps changing :(\[\Large s \sqrt3+\color{red}{27\sqrt{2^5}}+9\sqrt{3s}\]Is that red term correct? Or is the 2 supposed to be an `s` or something?

OpenStudy (anonymous):

\[\sqrt{27s ^{5}}\] This is the correct one... I'm not wearing my glasses and I'm tired (not the best mix)

zepdrix (zepdrix):

lol :)

zepdrix (zepdrix):

\[\Large s \sqrt3+\sqrt{27s^5}+9\sqrt{3s}\]Oh ok I see what you did there,\[\Large s \sqrt3+3s^2\sqrt{3s}+9\sqrt{3s}\]

zepdrix (zepdrix):

So I guess you're looking at it and thinking "Hey! they both have a sqrt(3s), can't we combine them??" But you want to look at the `entire` s value within each term. If I move things around, just a smidge! ...\[\Large s \sqrt3+3\sqrt3\color{royalblue}{s^2\sqrt{s}}+9\sqrt3\color{orangered}{\sqrt{s}}\]Can you see how the s's are quite different between the second and third term?

OpenStudy (anonymous):

Wouldn't it just stay outside of the radical?

zepdrix (zepdrix):

hmm? :o

zepdrix (zepdrix):

You could factor a sqrt(s) out of each term, yes.\[\Large \sqrt s\left(3\sqrt3 s^2+9\sqrt3\right)\]But I don't think that's what they had in mind when they asked you to `combine` them.

zepdrix (zepdrix):

You can't directly add them.

OpenStudy (anonymous):

I just got it! It can't be a like term if the coefficients aren't the same!

zepdrix (zepdrix):

The coefficients? D: ehh we don't really care about those :x It's all about the `s`'s!! :D

OpenStudy (anonymous):

Yeah but if the radicals are the same, it doesn't matter because the coefficients are not... They are not like terms and so I can't add them together.

zepdrix (zepdrix):

Ok ok ok :) I just want to make sure we agree on what a `coefficient` is. coefficient refers to the number in front of the variable. So if we had:\[\Large 5x+3x\]We `can` combine these terms despise the fact that the coefficients are different.\[\Large 8x\] What's stopping us from combining our `s`'s is that they are not like terms (as you said), different exponent on each variable. Just in case there was any confusion on that. c: